OpenStudy (anonymous):

At a school supply store, binders cost $5 and pencil pouches cost $2. In one day, 42 of these items are sold for total sales of $144. Which system represents this (b represents binders and p represents pencil pouches), and how many binders were sold?

OpenStudy (bibby):

I think I got this down now :D 42 = b + p 144 = 5b + 2p p = 42-b 144 = 5b + 2(42-b) 144 = 5b + 84 - 2b 60 = 3b ...

OpenStudy (anonymous):

haha thanks